# Higgs Inflation Model with Small Non-Minimal Coupling Constant

### v4

@@ 7 Discussion
In the discussion of the basic ideas and main results of the paper, it is worth concentrating on the role played by the combination of the presence of two volume measures in the TMSM with the flatness of the TMT-effective potential.
@@
To understand how Higgs inflation becomes possible at small $\xi$, we should start by recalling the special properties of the model that allow us to overcome the apparent contradiction between physics at inflationary-scale energies and the energies of accelerator experiments. The choice of a small value of the non-minimal coupling constant $\xi$ can be consistent with the CMB data only if the primordial model parameter $\lambda$ of the Higgs field selfcoupling is very small: as was shown, $\lambda=2.3\cdot 10^{-11}$ if $\xi=1/6$. But such a value of $\lambda$ is completely unsuitable for the implementation of the GWS theory. The resolution of this contradiction in the studied TMSM turns out to be possible only due to two volume measures in the primordial action. The main consequence of this is that $\lambda$ becomes a kind of running classical coupling constant, the value of which at energies near the Higgs field vacuum turns out to be approximately $10^{10}$ times larger than during the inflation stage. The very possibility of such a huge difference is the result of the fact that the value of the coefficient $b_{k}$ in the volume element $(b_{k}\sqrt{-g}-\Upsilon)d^{4}x$ is chosen extremely close to 1: $b_{k}=1+\mathcal{O}(10^{-5})$. The coefficient $b_{k}$ can be interpreted as a factor describing the relative contributions of the volume measure densities $\sqrt{-g}$ and $\Upsilon$ to the total volume element. Therefore, such a small deviation of $b_{k}$ from 1 may indicate the existence of a more fundamental theory in which the original balance between $\sqrt{-g}$ and $\Upsilon$ is broken by some weak effect. Thus, in TMSM the contradiction described above is resolved in a fairly natural way.
@@
Now let us return to the problem mentioned in the introductory section 1 that the initial conditions for inflation are unnatural. Such claims are widespread in the literature and are usually based on statistical estimates that the initial conditions (1.1) required for inflation to begin are exponentially rare. This argument became even more convincing when it turned out that models with plateau-like potentials are the most preferable in light of the Planck’s CMB results. As the authors of the paper [30] rightly noted, for a flat potential of height $\sim 10^{-10}M_{P}^{4}$ one would expect that initial values of the kinetic and gradient energy densities exceeding $\sim 10^{-10}M_{P}^{4}$, up to the Planck scale, are the most probable; and conversely, values of these energies satisfying the conditions (1.1) and necessary for the onset of inflation have a much lower probability. However, it is easy to see that in the context of TMSM this argument turns out to be completely inapplicable. First of all, given the K-essence type structure, we know that according to condition (6.8), the initial energy density $\rho_{in}$ turns out to be negative if $\frac{1}{2}\dot{\varphi}_{in}^{2}$ exceeds $0.8\cdot 10^{-10}M_{P}^{4}$. But one can also make a more fundamental argument that is in no way connected with the K-essence type structure. To do this, we recall that, as discussed in detail in section 5, in our Universe the ratio of two volume measures described by the scalar function $\zeta$ must be strictly positive. In section 5.2, analyzing the constraint (5.2) determining the values of $\zeta$, we found that for the beginning of the classical evolution of the Universe it is necessary that the initial values $\dot{\varphi}_{in}$ and $\nabla\varphi_{in}$ satisfy the upper bounds (5.9) and (5.10), respectively. From this analysis it follows that $\zeta$ becomes negative if we wish to choose initial values $\dot{\varphi}_{in}$ and $\nabla\varphi_{in}$ that exceed these upper bounds. And this, firstly, contradicts the conditions of applicability of the equations obtained from the principle of least action and, secondly, means changing the orientation of the spacetime manifold of our Universe to the opposite.
@@
The absence of the problem of initial conditions for Higgs inflation in the developed model is part of a more general result, namely, that inflation can be preceded by other, rather exotic initial stages of classical cosmological evolution. Since this result has fundamental theoretical significance, it is worth discussing its origin on a qualitative level. Since at $\varphi>6M_{P}$ the TMT-effective potential has a plateau of almost constant height, the possibility of the onset of cosmological evolution turns out to be independent of the initial value $\varphi_{in}>6M_{P}$. Therefore, from the pair of initial conditions $(\varphi_{in},\dot{\varphi}_{in})$ the possibility of the onset of evolution turns out to depend only on $\dot{\varphi}_{in}$. On the other hand, the TMT primordial action (2.1) leads to a TMT-effective action (3.4)-(3.10) having a K-essence-type structure: the kinetic part of the TMT-effective Lagrangian (3.5) is a quadratic form in $\dot{\varphi}^{2}$. This structure implies that the condition of positive initial total energy density, $\rho_{in}>0$, allows the initial TMT-effective kinetic energy density to be negative (although $\dot{\varphi}^{2}>0$). This makes it possible that the evolution of the Universe starts with an equation of state $w=\frac{p}{\rho}<-1$, i.e. proceeds according to the dynamics known in dark energy models as phantom. Other pathological initial states with $w>-1$ are also possible. Cosmological solutions describing the stages with pathological dynamics are of great interest and will be studied elsewhere. However, if the evolution starts with normal dynamics, then from the very beginning it proceeds in the inflationary regime.

### v5

@@ 8 Discussion
In the discussion of the basic ideas and main results of the paper, it is worth concentrating on the role played by the combination of the presence of two volume measures in the TMSM with the flatness of the TMT-effective potential. To understand how Higgs inflation becomes possible at small $\xi$, we should start by recalling the special properties of the model that allow us to overcome the apparent contradiction between physics at inflationary-scale energies and the energies of accelerator experiments. The choice of a small value of the non-minimal coupling constant $\xi$ can be consistent with the CMB data only if the primordial model parameter $\lambda$ of the Higgs field selfcoupling is very small: as was shown, $\lambda=2.3\cdot 10^{-11}$ if $\xi=1/6$. But such a value of $\lambda$ is completely unsuitable for the implementation of the GWS theory. The resolution of this contradiction in the studied TMSM turns out to be possible only due to two volume measures in the primordial action. The main consequence of this is that $\lambda$ becomes a kind of running classical coupling constant, the value of which at energies near the Higgs field vacuum turns out to be approximately $10^{10}$ times larger than during the inflation stage. The very possibility of such a huge difference is the result of the fact that the value of the coefficient $b_{k}$ in the volume element $(b_{k}\sqrt{-g}-\Upsilon)d^{4}x$ is chosen extremely close to $\zeta_{v}$: $b_{k}=\zeta_{v}+\mathcal{O}(10^{-5})$.
It would be natural to assume that the dynamics of the fields obtained from the primordial action of the TMSM must ensure that all total volume measures $(b_{i}\sqrt{-g}\pm\Upsilon)d^{4}x=(b_{i}\pm\zeta)\sqrt{-g}d^{4}x$ are positive throughout the evolution of the Universe. Direct verification made in this paper and in paper [22] shows that this requirement is met. The only exception is the total volume measure $(b_{p}\sqrt{-g}-\Upsilon)d^{4}x$ with the Higgs field mass term in action (2.3) (and, in the more general formulation, eq.(7.2)). When $\zeta$ increases during cosmological evolution from $\zeta\approx 0$ at the inflationary stage to $\zeta=\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}>\frac{1}{2}$ near vacuum, at some point in cosmic time $\zeta$ intersects the value $b_{p}$ and the total volume measure $(b_{p}\sqrt{-g}-\Upsilon)d^{4}x=(b_{p}-\zeta)\sqrt{-g}d^{4}x$ becomes negative. This pre-geometric effect (see the last paragraph in section A.1) is of particular interest because it manifests itself as the appearance of a negative squared mass of the Higgs field.
The choice of the dimensionless parameters $b_{k}$ in the bosonic sector, $b_{l}$, $l=e,\mu,\tau$, and $b_{q}$, $q=u,c,t$, in the fermionic sector, deserves a separate discussion. It is tempting to explain the closeness of all these parameters to $\zeta_{v}$ by a violation of some initial balance between $\sqrt{-g}$ and $\Upsilon$. However, as long as the mechanism responsible for the breaking of this type of symmetry remains completely unknown, we are forced to regard the choice of these parameters simply as an adjustment to achieve agreement with the measured quantities.
@@
Now let us return to the problem mentioned in the introductory section 1 that the initial conditions for inflation are unnatural. Such claims are widespread in the literature and are usually based on statistical estimates that the initial conditions (1.1) required for inflation to begin are exponentially rare. This argument became even more convincing when it turned out that models with plateau-like potentials are the most preferable in light of the Planck’s CMB results. As the authors of the paper [30] noted, for a flat potential of height $\sim 10^{-10}M_{P}^{4}$ one would expect that initial values of the kinetic and gradient energy densities exceeding $\sim 10^{-10}M_{P}^{4}$, up to the Planck scale, are the most probable; and conversely, values of these energies satisfying the conditions (1.1) and necessary for the onset of inflation have a much lower probability. However, as shown in this paper, this argument turns out to be completely inapplicable in the context of TMSM. First of all, given the K-essence type structure, we know (see eq.(6.8)) that for the initial energy density $\rho_{in}$ to be positive, the initial value $\frac{1}{2}\dot{\varphi}_{in}^{2}$ must satisfy the condition $\frac{1}{2}\dot{\varphi}_{in}^{2}\lesssim 0.78\cdot U_{eff}^{(tree)}(\varphi_{(in)})$. However, a more fundamental argument can be advanced, unrelated to the K-essence structure. In this regard, we recall that the choice made in this paper for the solution of the cosmological equations applicable to the description of our Universe imposes the requirement that the scalar function $\zeta$ be strictly positive. Violating this constraint would, firstly, contradict the conditions for the applicability of the equations derived from the principle of least action, and, secondly, would imply a change in the orientation of the space-time manifold of our Universe during its evolution. In section 5.2, by analyzing the constraint (5.2) determining the values of $\zeta$, we found that in order to satisfy the requirement $\zeta>0$ at the beginning of the classical evolution of the Universe, the initial values $\dot{\varphi}_{in}$ and $(\nabla\varphi)_{in}$ must satisfy the upper bounds (5.9) and (5.10), with which conditions (1.1) are satisfied. This allowed us to conclude that there is a wide range of model parameters for which the conditions (1.1) for the onset of inflation are satisfied.
@@
The absence of the problem of initial conditions for Higgs inflation in the developed model is part of a more general result, namely, that inflation can be preceded by other, rather exotic initial stages of classical cosmological evolution. This result is of fundamental theoretical importance, and its origin deserves a qualitative discussion. Since at $\varphi>6M_{P}$ the TMT-effective potential has a plateau of almost constant height, the possibility of the onset of cosmological evolution turns out to be independent of the initial value $\varphi_{in}>6M_{P}$. Therefore, from the pair of initial conditions $(\varphi_{in},\dot{\varphi}_{in})$ the possibility of the onset of evolution turns out to depend only on $\dot{\varphi}_{in}$. On the other hand, the TMT primordial action (2.1) leads to a TMT-effective action (3.5)-(3.11) having a K-essence-type structure: the TMT-effective kinetic energy density $K_{eff}=\Big(K_{1}-\frac{3}{2}\tilde{K}_{2}\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}\Big)\cdot\frac{1}{2}\dot{\varphi}^{2}$, eq.(6.6), is a quadratic form in $\dot{\varphi}^{2}$. This structure implies that the condition of positive initial total energy density, $\rho_{in}>0$, allows the TMT-effective kinetic energy density $K_{eff}$ to be negative at the initial $\dot{\varphi}_{in}$. This makes it possible that the evolution of the Universe starts with an equation of state $w=\frac{p}{\rho}<-1$, i.e. proceeds according to the dynamics known in dark energy models as phantom. Other pathological initial states with w>1w>-1 are also possible. However, as it was shown in section 6.2, if the evolution starts with normal dynamics, then from the very beginning it proceeds in the inflationary regime.
Finally, it’s worth noting that the description of fermion preheating in the TMSM model, presented in section 7.3, yields quantitative results consistent with those obtained in conventional models. Taken together with the other results obtained in this paper, this can be seen as further argument in favor of the TMSM over conventional models.
Appendix A Justification of the TMT approach developed in the paper
A.1 Why Two-Measure theory
To justify the structure of the primordial action in the Two-Measure Theory [^23] For more systematic and complete presentation of the TMT see recent review [69] presented in section 2, a clearly formulated new idea is required: all possible degrees of freedom contained in the primordial action should be considered as independent of each other, and all relationships between them appear as a result of applying the principle of least action. For brevity, we will call this approach to constructing a theory "the principle of maximum dynamic independence (PMDI)". The key to this formulation is the word "all" in the expression "all possible degrees of freedom", and we will now discuss what is meant by this.
Let us start with the Palatini formulation of gravity, in which the metric tensor and the (affine) connection are treated as independent variables. As is well known, in the absence of a non-minimal coupling of matter with curvature, the theory of gravity obtained by variation of the original action coincides with Einstein’s general relativity (GR). However, in the presence of, for example, a non-minimal coupling of a scalar field with a curvature scalar, the solution of the equations obtained by variation with respect to the affine connection leads to connection coefficients for which the covariant derivative of the metric tensor is non-zero. Violation of the metricity condition means that the space-time is non-Riemannian, and to formulate the theory in a Riemannian space-time, it is necessary to perform the corresponding Weyl transformation of the metric tensor. Typically, such a transformation is performed in the original action in such a way as to eliminate the non-minimal coupling of the scalar field with the curvature scalar, and the set of variables used in this case is called the Einstein frame.
Considering the formulation of Palatini’s gravity from the point of view of the applied mathematical apparatus, we inevitably come to the conclusion that we are dealing with a 4-dimensional differentiable manifold at the stages of equipping it with an affine connection and a metric structure (see, for example, [56]). The construction of the TMT can be considered as the next step in the implementation of the PMDI. Indeed, in an oriented 4-dimensional differentiable manifold, even before equipping it with an affine connection and a metric structure, one can define a volume measure as a 4-form, which can be represented as in eq.(1.2), $dV_{\Upsilon}=\Upsilon d^{4}x=4!d\varphi_{1}\wedge d\varphi_{2}\wedge d\varphi_{3}\wedge d\varphi_{4}$. After the metric structure is defined on the manifold, if the latter is orientable, $dV_{g}=\sqrt{-g}d^{4}x$ can also be chosen as a 4-form to describe the volume measure. The possibility of such a choice is based on the fact that under general coordinate transformations with a positive Jacobian (i.e. without changing the orientation of the 4-manifold), $\sqrt{-g}$ and $\Upsilon$ are transformed according to the same rule. The standard approach (see section 2.8 in ref.[56]) is to use only the volume element $dV_{g}$, by default assuming that the volume element $dV_{\Upsilon}$ is redundant in this situation. However, a closer look at the problem, when following PMDI, functions $\varphi_{a}$ are added to the set of degrees of freedom in action, shows that:
1. Only in empty space does the solution of the equations of the gravity model with two measures lead to the result $\Upsilon=const.\cdot\sqrt{-g}$, i.e. there is no fundamental difference between the volume measures $dV_{\Upsilon}$ and $dV_{g}$. However, even in empty space the gravity model with two measures can acquire new interesting properties compared to the gravity model with only one measure $dV_{g}$.
2. When the original action contains terms describing matter fields, and the volume measure $dV_{\Upsilon}$ is included in the action in addition to $dV_{g}$, then the scalar $\zeta\equiv\Upsilon/\sqrt{-g}$ turns out to be a function of the matter fields, which radically changes the dynamics of the fields.
The validity of the assertion 2 was shown in section 2, where the effect of two volume measures on the field dynamics was studied in detail. It remains to demonstrate the validity of the assertion 1, namely, that using both volume measures in empty space does indeed lead to Einstein’s gravity model, in which after solving the equations it is found that there is no fundamental difference between the volume measures $dV_{g}$ and $dV_{\Upsilon}$. This will be done in subsection A.2. When constructing models with two volume measures, we assume that the total volume measures have the form of linear combinations $(\alpha\sqrt{-g}+\beta\Upsilon)d^{4}x$, in which the constant dimensionless real coefficients $\alpha$, $\beta$ can be different depending in which term of the Lagrangian density they are present with in the primordial action. Bearing in mind that gravity in GR has a geometric nature, we will use the term pre-geometry for effects caused by the structure of the total volume measures used in the primordial action. In subsection A.3 we will discuss some additional very important properties of the pre-geometry.
A.2 Empty space
In examining the TMT model in empty space, we first consider the purely gravitational part of the primordial action, which in general can be chosen as $\int(\alpha\sqrt{-g}+\beta\Upsilon)\left(-\frac{M_{P}^{2}}{2}\right)g^{\mu\nu}R_{\mu\nu}(\Gamma)d^{4}x$, where $\alpha$ and $\beta$ are arbitrary parameters. The coefficient $\alpha$ can be extracted as a common factor from the parentheses and absorbed by redefining the original parameter $M_{P}$ (we keep the original notation for $M_{P}$). After this, by redefining the functions $\varphi_{a}$ by $(\frac{\beta}{\alpha})^{1/4}\varphi_{a}\to\varphi_{a}$, we reduce the purely gravitational part of the primordial action to the form $\int(\sqrt{-g}+\Upsilon)\left(-\frac{M_{P}^{2}}{2}\right)g^{\mu\nu}R_{\mu\nu}(\Gamma)d^{4}x$, which will be used in what follows. So, let us consider the TMT model in the empty space with the following primordial action
$S_{empty}=\int\Biggl[(\sqrt{-g}+\Upsilon)\left(-\frac{M_{P}^{2}}{2}\right)g^{\mu\nu}R_{\mu\nu}(\Gamma)-\sqrt{-g}V_{1}-\frac{\Upsilon^{2}}{\sqrt{-g}}V_{2}\Biggr]d^{4}x,$ (A.1)
The notation in (A.1) implying the use of Palatini’s formulation and the explanation of the presence of two vacuum-like terms were given in section 2.1. Here it is assumed that $V_{1}<0$ and $V_{2}<0$. The action takes into account the effect of pre-geometry and includes the following independent variables: the affine connection $\Gamma^{\lambda}_{\mu\nu}$, the metric tensor $g_{\mu\nu}$ and four scalar functions $\varphi_{a}$, (a=1…4), with the help of which the measure density $\Upsilon$ is constructed. Varying the action with respect to the functions $\varphi_{a}$, similar to what was done in section 2.2 (see eqs (2.6), (2.8), (2.10)), under condition $\Upsilon(x)\neq 0$ we obtain that the equality
$\frac{M_{P}^{2}}{2}g^{\mu\nu}R_{\mu\nu}(\Gamma)={\mathcal{M}}+2\zeta V_{2}$ (A.2)
must be satisfied, where $\zeta=\frac{\Upsilon}{\sqrt{-g}}$ and $\mathcal{M}$} is a constant of integration. Variation with respect to $g^{\mu\nu}$ yields the equation
$(1+\zeta)\left(-\frac{M_{P}^{2}}{2}\right)R_{\mu\nu}(\Gamma)-\frac{1}{2}g_{\mu\nu}\left(-\frac{\tilde{M}_{P}^{2}}{2}\right)g^{\alpha\beta}R_{\alpha\beta}(\Gamma)+\frac{1}{2}g_{\mu\nu}(V_{1}-\zeta^{2}V_{2})=0.$ (A.3)
The condition for compatibility of eqs.(A.2) and (A.3) is a constraint in the form of the following expression for a constant scalar $\zeta$
$\zeta=\frac{{\mathcal{M}}-2V_{1}}{{\mathcal{M}}-2V_{2}},$ (A.4)
As noted above in assertion 1, the latter means that the solution in empty space reveals that the volume measures $dV_{\Upsilon}$ and $dV_{g}$ differ only by a constant factor. From the constancy of $\zeta$ it follows that the variation of $\Gamma^{\lambda}_{\mu\nu}$ leads to an equation whose solution is the equality $\Gamma^{\lambda}_{\mu\nu}=\{^{\lambda}_{\mu\nu}\}$, where $\{^{\lambda}_{\mu\nu}\}$ are the Christoffel’s connection coefficients of the metric $g_{\mu\nu}$. This means that spacetime is Riemannian, there is no need for a Weyl transformation and $R_{\mu\nu}(\Gamma)=R_{\mu\nu}(g)$ is the Ricci tensor of the spacetime with the metric $g_{\mu\nu}$. Substituting the value of $\tilde{\zeta}$ into eq.(A.3), we obtain Einstein’s equations in empty space
$R_{\mu\nu}(g)=-\frac{1}{2M_{P}^{2}}\cdot\frac{{\mathcal{M}}^{2}-4V_{1}V_{2}}{{\mathcal{M}}-2V_{2}}g_{\mu\nu}.$ (A.5)
An important novelty of the solution of the model in empty space in TMT compared to the convensional theory is that by choosing the constant of integration $\mathcal{M}$}, we can implement a solution with any TMT-effective cosmological constant $\Lambda$. In particular, by choosing ${\mathcal{M}}=2\sqrt{V_{1}V_{2}}$, we obtain a solution with $\Lambda=0$ and
$\zeta|_{\Lambda=0}=\sqrt{\frac{V_{1}}{V_{2}}}.$ (A.6)
Note that in this case, the value of $\zeta|_{\Lambda=0}$ coincides, up to a very small correction, with the value $\zeta_{v}$ in the Higgs field vacuum found in section 2.3.
A.3 Adding matter fields
Continuing the justification for the choice of the primordial action (2.1), we take the action in empty space (A.1) as the gravitational term to which the matter terms must be added. As matter we consider the electroweak SM, and to describe it in the context of TMT we must take into account that in addition to rewriting the SM action in a general coordinate-invariant form, it is necessary to choose the volume measures with which the SM Lagrangian terms enter the primordial action. But before doing this, we must return to the mathematical issue discussed in section A.1 about the existence of two volume measures $dV_{g}$ and $dV_{\Upsilon}$ and take into account important differences between them. The density $\sqrt{-g}$ of the volume measure $dV_{g}$ is positive-definite and the volume measure $dV_{g}$ is defined if the manifold is oriented, [56]. But the density $\Upsilon$ of the volume measure $dV_{\Upsilon}$ is sign-indefinite and, moreover, the sign of the 4-form $dV_{\Upsilon}=4!d\varphi_{1}\wedge d\varphi_{2}\wedge d\varphi_{3}\wedge d\varphi_{4}$ determines the sign of the orientation of the manifold (for more details see, for example, [55]). Therefore, using linear combinations $(\alpha_{i}\sqrt{-g}+\beta_{i}\Upsilon)d^{4}x$ (’i’ lists the terms in the action) as a total volume measures in constructing the terms of the primordial action, we leave the sign of the space-time orientation undefined. And only as a result of choosing a solution to the equations of motion, the sign of $\Upsilon=\zeta\sqrt{-g}$ is fixed, and, therefore, the sign of the orientation. In essence, we are discovering the TMT-effect of spontaneous restoration of the sign of the space-time orientation. To understand how this effect occurs, let us consder the simplest example described by eqs. (A.4) and (A.5) (for more complicated case see eqs.(2.25)-(2.27) in section 2.3). To obtain a solution with $\Lambda=0$, we choose the integration constant ${\mathcal{M}}=2\sqrt{V_{1}V_{2}}$ and as a result found that $\zeta=\sqrt{\frac{V_{1}}{V_{2}}}>0$ and, therefore, $\Upsilon>0$. However, we could have chosen ${\mathcal{M}}=-2\sqrt{V_{1}V_{2}}$, which also ensures $\Lambda=0$. But with such a choice [^24] In the case studied in section 2.3, instead of the choice (2.26) for the integration constant $\mathcal{M}$} we could have chosen ${\mathcal{M}}=-2\sqrt{V_{1}V_{2}}(1+\delta)$. But with such a choice, the constraint (2.22) leads to $\zeta_{v}=-\sqrt{\frac{V_{1}}{V_{2}}}+\mathcal{O}\left(\frac{\lambda\sigma^{4}}{|V_{2}|}\right)$., the constraint (A.4) leads to $\zeta=-\sqrt{\frac{V_{1}}{V_{2}}}<0$ and, therefore, $\Upsilon<0$. Therefore, changing the sign of the integration constant $\mathcal{M}$}, we obtain solutions with opposite signs of the space-time orientation.
The conclusion that follows from the described TMT effect of spontaneous restoration of space-time orientation is so fundamental that it allows us to formulate in the most concentrated form the main difference between TMT and conventional field theories. In conventional field theories, it is assumed by default that the sign of space-time orientation can be chosen arbitrarily. But in TMT, as we have seen, one of the two possible signs of orientation of the space-time manifold of our Universe is fixed by the choice of the solution of cosmological equations. And since the sign of orientation coincides with the sign of the scalar function $\zeta$, which plays a decisive role in the dynamics of fields, the physical results predicted by TMT turn out to be different for different signs of the space-time orientation.
Since the primordial action does not imply the existence of a definite space-time orientation, and also because of the the sign-indefinitness of $\Upsilon$, when constructing the primordial action terms using linear combinations of $\sqrt{-g}$ and $\Upsilon$, it should be assumed that the coefficients in front of $\Upsilon$ in the total volume measures can be positive as well as negative. That is, assuming $\alpha_{i}>0$, $\beta_{i}>0$, the total volume measures can have the form $(\alpha_{i}\sqrt{-g}\pm\beta_{i}\Upsilon)d^{4}x$. This conclusion is a direct consequence of the fact that the structure of the primordial action takes into account pre-geometry in an essential way, and this is crucial for achieving the main results of this paper. and in paper [22]. Below we give a qualitative description of how this works. But before that, one more remark, more of a technical nature, needs to be made. Let $\tilde{L}_{i}$ denote the i-th term of the SM Lagrangian that enters the primordial action with volume measure $(\alpha_{i}\sqrt{-g}\pm\beta_{i}\Upsilon)d^{4}x$. Since the functions $\varphi_{a}$ have already been redefined to obtain the gravitational term (A.1) of the action, to eliminate the coefficient $\beta_{i}$, we can factor it out of the brackets and absorb it by redefining either the SM parameters or the SM fields in $\tilde{L}_{i}$. As a result, the i-th term of the Lagrangian density takes the form $(b_{i}\sqrt{-g}\pm\Upsilon)L_{i}d^{4}x$, where $b_{i}=\frac{\alpha_{i}}{\beta_{i}}$ and $L_{i}$ is obtained from the mentioned redefinitions in $\tilde{L}_{i}$. This is the form in which all the terms of the Two-Measure Standard Model are written in this paper and in paper [22].
We can now explain how the mathematical features of TMT described above are related to the possibility of obtaining the results of this paper and ref.[22]. The main role in this is played by the choice of signs in front of $\Upsilon$ in the volume measures $(b_{i}\sqrt{-g}\pm\Upsilon)d^{4}x$  : with a minus sign in the kinetic terms of all SM fields and with a plus sign in all other terms. In the framework of this paper, where gauge and fermion fields are completely excluded from consideration, the corresponding primordial action for the Higgs field is given by eq.(2.3). Due to this structure of the primordial action, during the evolution of the Universe, the TMT-effective selfcoupling constant of the Higgs field changes from the parameter value $\lambda\sim 10^{-11}$ in the inflationary epoch to the value $\lambda_{\sigma}$, eq.(2.32), related to the GWS theory. This occurs due to the change of $\zeta$ from a value close to zero during the inflationary epoch to the value $\zeta_{v}=\sqrt{\frac{V_{1}}{V_{2}}}$ near vacuum. In the latter case, this leads to the appearance in the denominator of the expression (2.32) for $\lambda_{\sigma}$ of a very small number $b_{k}-\zeta_{v}$, which compensates for the smallness of the parameter $\lambda$.
The only exception in the choice of signs in the volume measures is the minus sign in front of $\Upsilon$ in the volume measure of the Higgs field mass term, that allows to provide a mechanism for the change of the Higgs mass squared term during cosmological evolution from positive to negaive (see also the last paragraph in sec.4)
Appendix B Kinetic energy and $\zeta$ during inflation in the slow-roll regime
In this appendix we will analyze the function $\zeta(\varphi,X_{\varphi})$ on a flat segment of the TMT-effective potential $U_{eff}(\varphi)$, where $U_{eff}(\varphi)\approx 10^{-10}M_{P}^{4}$. The latter is realized at $\varphi\gtrsim 6M_{P}$, where the constraint that provides us with the function $\zeta(\varphi,X_{\varphi})$ is given by eq.(5.2), and for convenience we reproduce it here once again:
$\zeta(\varphi,X_{\varphi})=\frac{32q^{4}e^{-4\sqrt{\xi}\frac{\varphi}{M_{P}}}-\frac{\lambda}{8\xi^{2}}\delta_{p}+\frac{X_{\varphi}}{M_{P}^{4}}}{32q^{4}e^{-4\sqrt{\xi}\frac{\varphi}{M_{P}}}+\frac{\lambda}{8\xi^{2}}-\frac{X_{\varphi}}{M_{P}^{4}}},$ (B.1)
Let’s remember that $\lambda\approx 2.3\cdot 10^{-11}$,   $\xi=\frac{1}{6}$,   $q^{4}=3\cdot 10^{-10}$,   $b_{p}=\frac{1}{2}(1+\delta_{p})$,   $\delta_{p}\approx 10^{-8}$, that is $\frac{\lambda}{8\xi^{2}}\approx 10^{-10}$ and $32q^{4}e^{-4\sqrt{\xi}\frac{\varphi}{M_{P}}}\lesssim 5.3\cdot 10^{-13}$ for $\varphi\gtrsim 6M_{P}$. In the context of the FLRW Universe, we are limited to $X_{\varphi}$ as kinetic energy, that is, $X_{\varphi}=\frac{1}{2}\dot{\varphi}^{2}$. In sections 6.1 and 6.2 it was shown that the classical cosmological evolution starts with a normal dynamics only if the initial value $\dot{\varphi}_{in}$ satisfies the condition $\frac{1}{2}\dot{\varphi}_{in}^{2}<0.17\cdot U_{eff}^{(tree)}(\varphi_{in})$. Moreover, under this condition, cosmological evolution begins with inflation.
Taking into account that for $\varphi>6M_{p}$ with very high accuracy $K_{1}(\varphi)=1$, the cosmological eqs (6.24) and (6.25), describing inflation in the slow-roll approximation, are reduced to
$\frac{\dot{a}^{2}}{a^{2}}=\frac{1}{3M_{P}^{2}}U_{eff}^{(tree)}(\varphi),\qquad 3\frac{\dot{a}}{a}\dot{\varphi}=-U_{eff}^{(tree)\prime}(\varphi).$ (B.2)
From these equations it follows that in the slow-roll approximation the following equality is satisfied
$X_{\varphi}=\frac{1}{2}\Bigl(\frac{d\varphi}{dt}\Bigr)^{2}=\frac{M_{P}^{2}}{6U_{eff}}(U_{eff}^{(tree)^{\prime}})^{2}.$ (B.3)
From this we find that for $\varphi>6M_{P}$ the contribution of $X_{\varphi}$ to the constraint (B.1) is bounded from above according to the following estimate
$\frac{X_{\varphi}}{M_{P}^{4}}\approx 128\lambda\cdot e^{-2\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}<9\cdot 10^{-18}.$ (B.4)
Then, using the parameter values mentioned after eq.(B.1), we see that the first term in the numerator of eq.(B.1) is leading and is about $10^{5}$ times larger than the second and third terms. As is known from the graph of $\zeta(\varphi)$ with the same set of parameters, curve 3 in figure 2, $\zeta(\varphi)$ crosses zero at $\varphi\approx 14M_{P}$. We conclude that in the interval of $6M_{P}<\varphi<14M_{P}$, where $U_{eff}^{(tree)}(\varphi)$ is flat and the slow-roll approximation is valid, the scalar $\zeta$ can vary within the range
$0<\zeta\approx\frac{32q^{4}}{\lambda/8\xi^{2}}\cdot e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}\lesssim 5\cdot 10^{-3}.$ (B.5)